OpenStudy (anonymous):

integrate (x^2+8x+128)/(x^3-64)

OpenStudy (cggurumanjunath):

can you factorise (x^2+8x+128) ?

OpenStudy (anonymous):

I've got it mostly worked out but I keep getting the answer wrong

OpenStudy (anonymous):

I factored the bottom and did fraction decomp

OpenStudy (cggurumanjunath):

can you tell me what factors did you get for the numerator ?

OpenStudy (anonymous):

\[x ^{2}+8x+128=\frac{ A }{ x }+\frac{ B }{ (x+8) }+\frac{ C }{ (x-8) }\]

hartnn (hartnn):

that will be x^3 -64 x not x^3-64

OpenStudy (anonymous):

that's what it says, I didn't type it corrently

hartnn (hartnn):

x(x+8)(x-8) = x^3-64x but the Q has x^3-64

hartnn (hartnn):

oh, cool, the Q has x^3-64x ??

OpenStudy (anonymous):

yes, so it works out nicely

hartnn (hartnn):

yeah, what values u got for A,B,C

OpenStudy (anonymous):

A=2, B=-1 and C=0?

hartnn (hartnn):

C = 0 ? no....how ?

OpenStudy (anonymous):

lol, that's where I figured I was messing up

OpenStudy (anonymous):

\[A+B+C=1\] \[-8B+8C=8\] 64A=128

OpenStudy (anonymous):

A=2

OpenStudy (anonymous):

\[B+C=-1\] -B+8C=8

OpenStudy (anonymous):

-8B

hartnn (hartnn):

to get C, you need to put x=8 in x^2+8x+128 = A (x+8)(x-8) + Bx(x-8) +C (x(x+8)) got that^ ???

OpenStudy (anonymous):

yeah, I was solving it differently but your way is probably easier

hartnn (hartnn):

then try it :)

OpenStudy (anonymous):

especially since mine didn't work

OpenStudy (anonymous):

C=2

hartnn (hartnn):

+2 or -2 ?

OpenStudy (anonymous):

I know it should be -2 but I'm not sure where the negative came from

hartnn (hartnn):

yeah, its +2 only

OpenStudy (anonymous):

so it is positive?

hartnn (hartnn):

oh, it should be +2 only

hartnn (hartnn):

what about A and B ? to get A, you need to put x=0 in x^2+8x+128 = A (x+8)(x-8) + Bx(x-8) +C (x(x+8)) to get B, you need to put x=-8 in x^2+8x+128 = A (x+8)(x-8) + Bx(x-8) +C (x(x+8))

hartnn (hartnn):

C= +2 is correct, if you had any doubts...

OpenStudy (anonymous):

So A and C are 2?

hartnn (hartnn):

A is not 2

hartnn (hartnn):

128 = A (-64)

OpenStudy (anonymous):

ohhhhhh

hartnn (hartnn):

so, B= ?

OpenStudy (anonymous):

So B is +1, not -1

OpenStudy (anonymous):

Awesome, I got the answer now

hartnn (hartnn):

yup u got it all

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome :)